the subject group head of shs mathematics department announced that the mean percentage score of grade 11 learning in the 3rd quarter grade in statistic and probability was it in 9 in the standard deviation was 12 one learned who believed that the mean percentage score was less than this randomly selected 34 learner in computed their mean scores obtained a mean percentage score of 85 at 0.01 level of significance test the learners belief

1
Expert's answer
2022-05-25T09:32:07-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge89"

"H_1:\mu<89"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a left-tailed test is "z_c =-2.3263."

The rejection region for this left-tailed test is "R = \{z:z<-2.3263\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{85-89}{12/\sqrt{34}}\approx-1.9437"

Since it is observed that "z|=-1.9437>-2.3263=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(z<-1.9437)= 0.025966," and since "p= 0.025966>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 89, at the "\alpha = 0.01" significance level.